The weights of a random sample of cereal boxes that are supposed to weigh 1 pound are given below. Estimate the standard deviation of the entire population with 98.7% confidence. Assume the population is normally distributed.

0.97, 1.05, 0.98, 1.06, 1.03, 0.99, 0.99, 1.01
Answer #1

Here s = 0.0334 and n = 8
df = 8 - 1 = 7

α = 1 - 0.987 = 0.013
The critical values for α = 0.013 and df = 7 are Χ^2(1-α/2,n-1) = 1.076 and Χ^2(α/2,n-1) = 19.601

CI = (sqrt(7*0.0334^2/19.601) , sqrt(7*0.0334^2/1.076))
CI = (0.02 , 0.0852)